OpenStudy (smurfy14):

Describe the sequence of transformations from the parent to the given function: f(x)=3(x-1)^2+2 help!!

OpenStudy (anonymous):

i guess the "parent" is \[g(x)=x^2\]

OpenStudy (anonymous):

\[g(x-1) = (x-1)^2 \] shifts the graph right 1

OpenStudy (anonymous):

\[3g(x-1)=3(x-1)^2 \] makes it narrower \[3g(x-1)^2+2=3(x-1)^2+2\] raised it two units up

OpenStudy (smurfy14):

so i'd put it became narrower and moved one unit to the right and two units up?

OpenStudy (smurfy14):

and instead of "it" would i say parabola?